# A Multimodal Dataset for Visually Grounded Ambiguity in Machine Translation

### 03 May 2026 v1

@@ A.2 VIDA Subset Statistics
Subset Ambiguity Focused Size Avg. Length (words) Avg. Ambi. Terms
VIDA-Base Word-Level 1,932 11.12 1.7826
VIDA-Sent Sentence-Level 312 6.00 1.00
VIDA-CollN Word-Level 256 10.08 1.20
Table 3: Statistical summary of VIDA subsets
Appendix B Additional Validation and Judge Details
B.1 Visual Dependence Ablations
@@
B.2 Judge Reliability
We further validate the reliability of the Judge from two complementary perspectives. First, we perform Cross-Model Judge Verification by re-evaluating the exact same translation outputs from Qwen2.5-VL-7B (CoT-SFT) on All-Test with a second Judge based on LLaMA3.1-8B. As shown in Table 6, the scores remain highly consistent across judge backbones. Second, we measure Human-Judge Alignment on a randomly sampled 20% subset of VIDA (500 instances) and compute Cohen’s Kappa between expert human judgments and the Qwen-based Judge. Table 7 shows strong agreement across all subsets, supporting that the Judge captures semantic disambiguation resolution rather than model-specific preferences.
@@ B.2 Judge Reliability
B.3 Judge Training Details
@@
We fine-tune the Judge for 10 epochs with learning rate $1\times 10^{-5}$ and batch size 8. LoRA is used with rank 8, alpha 32, and dropout 0.1. The final training set is constructed with a positive-to-negative ratio of 2:1. This contrastive setup encourages the Judge to focus on disambiguation correctness rather than superficial lexical overlap, which is crucial for handling valid paraphrases and lexical variation in open-ended translation.
@@
Appendix C Chain-of-Thought Supervised Fine-Tuning
@@ Appendix D Qualitative Analysis
The VIDA-CollN example (left of Figure 3) illustrates the collective noun ambiguity, which the source sentence contains the ambiguous noun "object", which requires a concrete translation ("paddle"). The SFT model, without reasoning, outputs the literal "object," which fails to capture the intended meaning. In contrast, the CoT-SFT shows that the model first generates an initial translation ( 物体 ), maintaining the literal meaning. During the ambiguity check, the model detects that "object" is ambiguous. In the subsequent visual disambiguation step, it grounds the word to the image and identifies that the woman is holding a paddle. Finally, in localized refinement, the model updates the translation to "paddle", producing the correct disambiguated output.
The VIDA-Sent example (middle of Figure 3) demonstrates sentence-level ambiguity where an idiomatic expression could be misunderstood literally. The phrase "got a green thumb" could be interpreted literally or idiomatically. The SFT model again produces a literal output of the color of thumb in Chinese. In contrast, the CoT-SFT first provides a literal initial translation ( 绿色的手 ). Through visual disambiguation, it recognizes from the image that the woman is gardening, and therefore refines the output to "Gardening expert", correctly capturing the idiomatic meaning.
@@
The right panel of Figure 3 illustrates an overthinking case. The phrase "iPod’s touch" should be interpreted as "iPod-like touch screen". The model first provides a reasonable image description and recognizes the intended interpretation during ambiguity checking. However, in the later disambiguation step, it over-interprets the phrase by incorrectly linking it to "someone physically touching" mentioned in the grounding step, rather than the relevant cue about the product feature. As a result, the model revises an initially adequate interpretation into an incorrect final translation. This example suggests that excessive reasoning can override correct early hypotheses and partially explains the performance drop observed for CoT-SFT on in-distribution data.

### 26 May 2026 v2

@@ Appendix C Dataset Statistics
Criterion Cohen’s $\kappa$
Disambiguation 0.8033
Fluency 0.7938
Preservation 0.8102
Average 0.8024
Table 5: Inter-annotator agreement for human verification.
@@
C.1 Definition and Detection of Ambiguous Terms
We define an ambiguous term as a word or phrase whose translation admits multiple valid candidates when given only the pure text context. Crucially, its true meaning can only be uniquely determined by referencing the paired visual evidence. Such terms are detected during Stage 1 of our curation pipeline (section 3): Qwen-Max and DeepSeek-v3 independently evaluate the text-only source for translation uncertainty, and an instance is retained only when both models reach a consensus that ambiguity exists. The two models then extract the specific ambiguous span and produce an Ambiguity Rationale that describes the alternative interpretations and identifies the visual cue required for disambiguation. Concrete examples are illustrated in Figure 3.
C.2 Pipeline Filtering Statistics
The large reduction from the raw candidate pool to the final benchmark reflects the strict quality control in our curation pipeline. We begin with 26,452 raw image–text pairs, retain 14,993 after image–text matching and text normalization, and then keep only 2,500 instances after ambiguity detection and dual-LLM verification. The two models show strong bidirectional agreement on the ambiguity decision: 94.38% of ambiguous candidates flagged by Qwen-Max are independently confirmed by DeepSeek-v3, and 87.50% in the reverse direction, indicating that the detected ambiguities are robust rather than model-specific artifacts. Subsequent stages refine translations and annotations without discarding additional instances. All 2,500 retained instances are then human-verified by two native Chinese annotators, with substantial inter-annotator agreement across the three review criteria (average Cohen’s $\kappa=0.8024$; Table 5). The filtering funnel is summarized in Table 4.
C.3 VIDA Subset Statistics
The rigorous pipeline outlined in section 3 results in the construction of VIDA (Visually-Dependent Ambiguity), comprising 2,500 visually dependent instances across three subsets: VIDA-Base contains 1,932 samples curated from 3AM and focuses primarily on word-level ambiguities; VIDA-CollN contains 256 collective-noun cases where visual context concretizes an otherwise abstract group; and VIDA-Sent contains 312 samples adapted from MMA with sentence-level semantic ambiguities. A complete statistical summary is provided in Table 6.
Subset Ambiguity Size Avg. Len. Avg. Ambi.
VIDA-Base Word-level 1,932 11.12 1.78
VIDA-Sent Sentence-level 312 6.00 1.00
VIDA-CollN Word-level 256 10.08 1.20
Table 6: Statistical summary of VIDA subsets.
C.4 Visual Dependence
Appendix D Judge Reliability and Training Details
D.1 Judge Reliability
We further validate the reliability of the Judge from two complementary perspectives. First, we perform Cross-Model Judge Verification by re-evaluating the exact same translation outputs from Qwen2.5-VL-7B (CoT-SFT) on All-Test with a second Judge based on LLaMA3.1-8B. As shown in Table 9, the scores remain highly consistent across judge backbones. Second, we measure Human-Judge Alignment on a randomly sampled 20% subset of VIDA (500 instances) and compute Cohen’s Kappa between expert human judgments and the Qwen-based Judge. Table 10 shows strong agreement across all subsets, indicating that the Judge captures semantic disambiguation resolution rather than model-specific preferences.
@@ D.1 Judge Reliability
D.2 Judge Training Details
@@
We fine-tune the Judge for 10 epochs with learning rate $1\times 10^{-5}$ and batch size 8. LoRA is used with rank 8, alpha 32, and dropout 0.1. The final training set is constructed with a positive-to-negative ratio of 2:1. This contrastive setup encourages the Judge to focus on disambiguation correctness rather than superficial lexical overlap, which is crucial for handling valid paraphrases and lexical variation in open-ended translation. All experiments were conducted on one NVIDIA A100 GPU.
@@
Appendix E Chain-of-Thought Supervised Fine-Tuning
@@ Appendix F Qualitative Analysis
The VIDA-CollN example (left of Figure 3) illustrates the collective noun ambiguity: the source sentence contains the ambiguous noun "object", which requires a concrete translation ("paddle"). The SFT model, without reasoning, outputs the literal "object," which fails to capture the intended meaning. In contrast, the CoT-SFT shows that the model first generates an initial translation ( 物体 ), maintaining the literal meaning. During the ambiguity check, the model detects that "object" is ambiguous. In the subsequent visual disambiguation step, it grounds the word to the image and identifies that the woman is holding a paddle. Finally, in localized refinement, the model updates the translation to "paddle", producing the correct disambiguated output.
The VIDA-Sent example (middle of Figure 3) demonstrates sentence-level ambiguity where an idiomatic expression could be misunderstood literally. The phrase "got a green thumb" could be interpreted literally or idiomatically. The SFT model again produces a literal Chinese rendering of the thumb color. In contrast, the CoT-SFT first provides a literal initial translation ( 绿色的手 ). Through visual disambiguation, it recognizes from the image that the woman is gardening, and therefore refines the output to "gardening expert", correctly capturing the idiomatic meaning.
@@
The right panel of Figure 3 illustrates an overthinking case. The phrase "iPod’s touch" should be interpreted as "iPod-like touch screen". The model first provides a reasonable image description and recognizes the intended interpretation during ambiguity checking. However, in the later disambiguation step, it over-interprets the phrase by incorrectly linking it to "someone physically touching" mentioned in the grounding step, rather than the relevant cue about the product feature. As a result, the model revises an initially adequate interpretation into an incorrect final translation. This example suggests that excessive reasoning can override correct early hypotheses and partially explains the performance drop observed for CoT-SFT on in-domain data.
Appendix G Prompts
This section lists the prompts used in VIDA construction and disambiguation evaluation.
Prompt 1: Stage 1 Ambiguity Detection Prompt 2: Stage 1 Consensus Merging Prompt 3: Stage 2 Disambiguation Translation Prompt 4: Disambiguation Judge